OpenStudy (anonymous):

Calculus III Questions: 1. Let P be the point (1, 2, 3, 4) and Q the point (4, 3, 2, 1). Let A be the vector <1, 1, 1, 1>. Let L be the line passing through P and parallel to A. (a) Given a point X on the line L compute the distance between Q and X (as a function of parameter t). (b) Show that there is precisely one point X0 on the line such that this distance achieves a minimum, and that this minimum is 2sqrt(5). (c) Show that X0-Q is perpendicular to the line.

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

It's a textbook question in the chapter on differentiation of vectors. So I know I have to get the derivative somewhere, I just don't know how to approach the problem.

OpenStudy (phi):

This is a bit complicated. First, the equation of the line,using vectors will be L = P + tA (Caps will mean a vector, lower case a scalar) (a) distance between points Q and X (X is on the line) is | Q-X| = sqrt |Q-X|^2 = sqrt ( (Q-X)' (Q-X) ) where ' means transpose. As the expression is unwieldy, let's find the length squared. Because X is on the line, X= P + tA sub into the expression for magnitude squared (Q-P -tA)' (Q-P -tA) expand: distance squared = Q'Q-P'Q-tQ'A-P'Q+P'P+tP'A-tA'Q+tA'P+t^2A'A distance squared = Q'Q -2P'Q +P'P -2tQ'A +2t P'A + t^2 A'A (using the fact that P'Q= Q'P and P'A= A'P) (b) take the derivative with respect to t and set to zero, solve for t X0= P+ t0 A where t0 is the solution just found. find the length of vector Q-X0 Use |Q-X0|^2 =(Q-X0)' (Q-X0) = 20, and sqrt(20)= 2sqrt(5) (c) Let Y = X0-Q, and show that Y'A = 0 (where A is the slope of the line)

OpenStudy (phi):

I left a lot of the work unspecified. If you get stuck, ask more questions.

OpenStudy (anonymous):

when you write ' means transpose, do you mean ' = derivative?

OpenStudy (anonymous):

i've never seen the term "transpose"

OpenStudy (phi):

no, I mean transpose of a vector. Normally you use T as an exponent

OpenStudy (phi):

Do you know "dot product" of two vectors?

OpenStudy (anonymous):

of course

OpenStudy (phi):

when you multiply vectors , they have to have the same shape. In this case, P'Q (for example) means P dot Q

OpenStudy (phi):

How would you multiply P times Q?

OpenStudy (anonymous):

P*Q = <4, 6, 6, 4> * = dot product

OpenStudy (phi):

OK, that will work. Treat all the A'B as A dot B. Hopefully it makes sense

OpenStudy (anonymous):

(a) makes sense for (b) why does (Q-X0)*(Q-X0) = 20? where is 20 from

OpenStudy (anonymous):

waaiit. nvm

OpenStudy (anonymous):

(b) makes sense as well

OpenStudy (phi):

(b) is where you have to do the work. That is the answer

OpenStudy (anonymous):

and for c it's just use of the dot product because dot product = 0 means perpendicular.

OpenStudy (anonymous):

how do you get A - the slope of the line for (c)

OpenStudy (phi):

yes. It isn't too hard to finish off (b) and (c) (I hope!)

OpenStudy (anonymous):

rather, what equation am I getting the derivative of for (c)?

OpenStudy (phi):

Let A be the vector <1, 1, 1, 1>. Let L be the line passing through P and parallel to A.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

cool!

OpenStudy (phi):

no derivative for (a) or (c).

OpenStudy (anonymous):

do you have time to answer another (less long) calc III question? it's hard to get good answerers for calc 3 stuff

OpenStudy (phi):

Is it posted? If not, post it, because this one is long.

OpenStudy (anonymous):

okay i'll make it a new post